OpenStudy (anonymous):

Standard form! help! (5,4)(-8,10) Only use integer coefficients. Write the standard form of the equation of the line these points go through.

OpenStudy (anonymous):

-6/13

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

Yes, that's your slope: \(\Large m=-\frac{6}{13} \) So the equation in point-slope form is: \(\Large y-4=-\frac{6}{13}(x-5) \) Clearing the fraction you can multiply by 13: \(\Large 13y-13\cdot4=-6(x-5) \) \(\Large 13y-52=-6x+30 \) So getting this in standard form Ax + By = C, \(\Large \color {red}{6x}+13y-52\color {blue}{+52}=-6x\color {red}{+6x}+30\color {blue}{+52} \) \(\Large 6x+13y=82 \)